# Energy-Efficient Port Selection and Beamforming Design for Integrated Data and Energy Transfer Assisted by Fluid Antennas

### v1

@@ Abstract
Integrated data and energy transfer (IDET) is considered as a key enabler of 6G, as it can provide both wireless energy transfer (WET) and wireless data transfer (WDT) services towards low power devices. Thanks to the extra degree of freedom provided by fluid antenna (FA), incorporating FA into IDET systems presents a promising approach to enhance energy efficiency performance. This paper investigates a FA assisted IDET system, where the transmitter is equipped with multiple FAs and transmits wireless signals to the data receiver (DR) and the energy receiver (ER), which are both equipped with a single traditional antenna. The switching delay and energy consumption induced by port selection are taken into account in IDET system for the first time. We aim to obtain the optimal beamforming vector and the port selection strategy at the transmitter, in order to maximize the short-term and long-term WET efficiency, respectively. The instant sub-optimal solution is obtained by alternatively optimizing the beamforming vector and port selection in each transmission frame, while a novel constrained soft actor critic (C-SAC) algorithm is proposed to find the feasible policy of port selection from the long-term perspective. Simulation results demonstrate that our scheme is able to achieve greater gain in terms of both the short-term and long-term WET efficiency compared to other benchmarks, while not degrading WDT performance.
@@ I-A Background
Physical design of FA mainly focuses on liquid-based antennas, reconfigurable pixel-based antennas and stepper motor based antennas [4]. The liquid based FA consists of liquid radiating elements (e.g., mercury, ga, galinstan, etc.) enclosed within a dielectric holder and controlled by a dedicated microelectromechanical system (MEMS) [5]. The reconfigurable pixel based FA is composed of an array of metasurface elements or reconfigurable sub-wavelength structures that can dynamically modify their electromagnetic properties [6][7] and the stepper motor based FA is known as the movable antenna (MA)[8] [9]. Compared to the high hardware costs of pixel-based antennas, which stem from a large number of fixed antennas, liquid-based FAs and MAs are more cost-effective as they require fewer antenna elements and RF chains. However, liquid-based FAs and MAs incur delay and energy consumption during the movement, which may limit their use in some communication environments. Therefore, evaluating the impact of FA movement is crucial for achieving higher energy efficiency and spectral efficiency of IDET systems.
@@ I-B Related works
In order to establish an energy-efficient interconnected network, IDET leverages RF signals to supply energy for low-power wireless devices. Plenty of works have been dedicated to the transceiver design of IDET system, aiming to improve the performance trade-off between WDT and WET. For instance, Kwon et al. [10] developed a transceiver design for a mmWave IDET system, including the transmit power allocation, beamforming and receive power splitting. Zhang et al. [11] considered a practical IDET scenario, where the data receivers (DRs) and energy receivers (ERs) were located in the far- and near-field regions of the extremely large-scale array base station (BS), respectively. The weight sum-power harvested at ERs was maximized by jointly optimizing the beamforming vectors and transmit power under the constraints of maximum sum-rate. With the assistance of reconfigurable intelligent surface (RIS), the performance of IDET system can be further improved [12, 13, 14]. Wu et al. [12] studied a sets of RIS assisted IDET system, where the beamforming vectors and reflect phase shifts were jointly optimized to minimize the transmit power, subject to the quality-of-service (QoS) constraints for all users. Yaswanth et al. [13] jointly optimized the beamforming matrices at both base station and RIS to minimize the transmit power while satisfying the minimum DR and ER constraints, which contributed to an energy efficient transmission scheme.
@@ I-C Motivation and Contributions
Although previous works, including our own work [23] have shown that deploying FA in IDET systems can achieve outstanding performance than traditional antenna, they have not considered the impact of switching delay and energy consumption associated with antenna movement, which may decrease the throughput and energy efficiency of IDET system. In IDET system, the WDT performance and WET performance can be improved by moving FAs into strategic positions, where the signal strength is maximized, or interference is minimized, or a trade-off between both factors is achieved. However, the movement of FAs introduces both delay and energy consumption, which reduces the effective transmission time within each frame, thereby lowering the achievable throughput of DR and harvested energy of ER. Therefore, it is crucial to balance the channel gain brought from the FAs and the energy efficiency loss caused by their movement. Motivated by this, we consider the impact of delay and energy consumption of FA movement on the IDET performance from both short-term perspective and long-term perspective, aiming to enhance IDET efficiency and provide some new insights. The contributions are summarized as follows:
@@
  • The WET efficiency is maximized by optimizing the beamforming vector and port selection from short-term perspective and long-term perspective, while ensuring the harvested energy constraint and achievable throughput constraint. In short-term optimization, we transform the original fraction problem into linear problem by using Dinkebach’s algorithm, while the port selection and beamforming vector are then optimized alternatively by using feasible point pursuit-successive convex approximation (FPP-SAC) and semidefinite relaxation (SDR).
@@
Notations: In this paper, boldface upper letters, boldface lower letters, and lower letters denote matrices, vectors, and scalars, respectively. $\mathbb{C}^{x\times y}$ and $\mathbb{R}^{x\times y}$ represent the set of complex matrices and real matrices with the dimension of $x\times y$, respectively. $|x|$ represents the amplitude of a complex number $x$. $||\boldsymbol{x}||$ and $||\boldsymbol{x}||_{1}$ denote the 2-norm and 1-norm of a vector $\boldsymbol{x}$, respectively. The notation $(\cdot)^{T}$ and $(\cdot)^{H}$ refer to the transpose and the conjugate transpose of a vector or matrix, while $\mathbf{E}$ represents the expectation operator.
@@ II System Model
As shown in Fig. 1, we consider a multiple FAs assisted IDET system, which is comprised of a transmitter, a DR and an ER. The transmitter is equipped with $N_{t}$ FAs, each of which has the same size and parameters without loss of generality. The DR and ER are both equipped with a single traditional antenna due to the limited hardware space. We assume that the transmit antennas are far apart in space and the channels of different antennas are fully independent, while the channels among different ports in the same antenna are strongly spatially correlated. The set of FAs and FA’s port are denoted by $\mathcal{N}_{t}=\left\{1,\cdots,N_{t}\right\}$ and $\mathcal{N}=\left\{1,\cdots,N\right\}$, respectively.
@@ II-B Wireless channel model
On the other hand, the channels of different transmission frames are temporally correlated, which can be expressed as
$\boldsymbol{h}_{i}(t)=\rho\boldsymbol{h}_{i}(t-1)+\sqrt{\frac{1-\rho^{2}}{PL_{h,i}}}\boldsymbol{\delta},$
$\boldsymbol{g}_{i}(t)=\rho\boldsymbol{g}_{i}(t-1)+\sqrt{\frac{1-\rho^{2}}{PL_{g,i}}}\boldsymbol{\delta}$ (5)
where $\rho$ is the temporal correlation factor and $\boldsymbol{\delta}\sim\mathcal{CN}(0,\boldsymbol{I}_{N})$ is the random difference with an identity matrix $\boldsymbol{I}_{N}$.
@@ II-C Signal model
where $\eta\in(0,1]$ represents the energy conversion efficiency111 In general, considering a nonlinear energy harvest model is more practical, but the received power at ER is much small due to the signal attenuation, enabling the EH circuit works in linear region of the nonlinear model [27]. Thus the linear EH model can be used to approximate the nonlinear model.. Assuming that the power actuating the liquid metal moving is a constant as $P_{u}$, the total energy consumption in coherence time is expressed as
$E_{c}(t)$ $=(T-\tau_{s}(t))||\boldsymbol{w}(t)||^{2}+\sum_{i=1}^{N_{t}}\tau_{i}(t)P_{u}+TP_{c},$ (11)
@@
$=$ $\frac{\eta(T-\tau_{s}(t))|\boldsymbol{g}_{s}(t)^{H}\boldsymbol{w}(t)|^{2}}{(T-\tau_{s}(t))||\boldsymbol{w}(t)||^{2}+\sum_{i=1}^{N_{t}}\tau_{i}(t)P_{u}+TP_{c}}.$ (12)
In this paper, we aim to investigate both short-term and long-term WET efficiency optimization for the FA assisted IDET system:
@@
  • Conversely, the long-term optimization framework adopts a holistic perspective by considering temporal correlations of channels and state evolution across multiple transmission frames, which can achieve more stable long-term WET efficiency. Nevertheless, the long-term optimization based DRL algorithms requires prior experience and the training is time-consuming.
Both schemes can provide valuable insights for the design of the FA assisted IDET system.
@@ III-A Problem Formulation
In this section, we optimize the beamforming vector and port selection to maximize the WET efficiency in each transmission frame by only exploiting the current channel state information (CSI). For the sake of simplicity, we omit the time index $t$ in the expression. Therefore, the optimization problem can be formulates as
[…]
@@
where $\boldsymbol{G}=\mathcal{M}\left\{\boldsymbol{g}_{i},i\in\mathcal{N}_{t}\right\}$ and $\boldsymbol{D}=\mathcal{M}\left\{\boldsymbol{d}_{i},i\in\mathcal{N}_{t}\right\}$. Note that the objective function is a complicated fraction and the beamforming vector and the port activation indicators are intricately coupled, thus (P2) is non-convex optimization problem and is hard to solve directly. According to [28], Dinkebach’s algorithm is applied to transform the fractional objective function to linear function. By relaxing the binary constraints and denoting $\boldsymbol{C}=\mathcal{M}\left\{\boldsymbol{1}_{i}^{N},i\in\mathcal{N}_{t}\right\}$, (P2) can be reformulated as
[…]
@@
$\hat{\boldsymbol{a}}\in[0,1]^{NN_{t}}$ (17g)
with a auxiliary variable $\beta$, which is iteratively updated by
$\beta^{n}=\frac{E_{h}^{n}}{E_{c}^{n}}$ (18)
where $n$ is the iteration index. For a fixed parameter $\beta$, (P3) is still a non-convex problem, thus we develop an AO algorithm to tackle it by solving $\hat{\boldsymbol{a}}$, $\tau_{s}$ and $\boldsymbol{w}$ iteratively.
@@ III-B Port selection optimization
$\eta|\hat{\boldsymbol{a}}^{T}\boldsymbol{G}\boldsymbol{w}|^{2}-\beta\left(||\boldsymbol{w}||^{2}+\frac{||\hat{\boldsymbol{a}}^{T}\boldsymbol{D}||_{1}P_{u}+TP_{c}}{(T-\tau_{s})}\right)$ (19)
$=\eta|\hat{\boldsymbol{a}}^{T}\boldsymbol{G}\boldsymbol{w}|^{2}-\beta\left(||\boldsymbol{w}||^{2}+\frac{||\hat{\boldsymbol{a}}^{T}\sqrt{\boldsymbol{D}}||_{2}^{2}P_{u}+TP_{c}}{(T-\tau_{s})}\right)$
@@
(P4) contains non-convex terms $|\hat{\boldsymbol{a}}^{T}\boldsymbol{G}\boldsymbol{w}|^{2}$ and $|\hat{\boldsymbol{a}}^{T}\boldsymbol{H}\boldsymbol{w}|^{2}$ in both the objective function (20) and constraints (20a)-(20b), while maintaining convexity in all other terms with respect to optimization variables $\hat{\boldsymbol{a}}$ and $\tau_{s}$. Although semidefinite relaxation (SDR) approach can be used to linearize the non-convex components, its application would induce significant dimensionality expansion of the optimization variables, rendering it computationally prohibitive for FA-assisted IDET system with large port number $N$. The FPP-SCA algorithm [29] is proposed for approximately solving general quadratic constraint quadratic problems (QCQPs), while a feasible point is guaranteed by adding slack variables. For any $\hat{\boldsymbol{a}}$,$\hat{\boldsymbol{x}}$, we have $(\hat{\boldsymbol{a}}-\hat{\boldsymbol{x}})^{T}\boldsymbol{S}(\hat{\boldsymbol{a}}-\hat{\boldsymbol{x}})\geq 0$, where $\boldsymbol{S}=\eta\boldsymbol{G}\boldsymbol{w}\boldsymbol{w}^{H}\boldsymbol{G}^{H}$. Expanding the right-hand side of the inequality, we obtain
@@
$-\beta\left(\frac{||\hat{\boldsymbol{a}}^{T}\sqrt{\boldsymbol{D}}||_{2}^{2}P_{u}+TP_{c}}{(T-\tau_{s})}\right),$ (22)
$s.t.$ $2Re\left\{\hat{\boldsymbol{a}}^{T}\boldsymbol{S}\hat{\boldsymbol{x}}\right\}-\hat{\boldsymbol{x}}^{T}\boldsymbol{S}\hat{\boldsymbol{x}}+s_{1}\geq\frac{E_{th}}{T-\tau_{s}},$ (22a)
@@
where $p$ is slack penalty, $s_{1}$ and $s_{2}$ are slack variables, $\boldsymbol{F}=\boldsymbol{H}\boldsymbol{w}\boldsymbol{w}^{H}\boldsymbol{H}^{H}$. Given an initial point $\hat{\boldsymbol{x}}$, (P5) is a convex optimization problem and can be optimally solved with the assistance of convex optimization tools.

### v2

@@ Abstract
Integrated data and energy transfer (IDET) is considered as a key enabler of 6G, as it can provide both wireless energy transfer (WET) and wireless data transfer (WDT) services towards low power devices. Thanks to the extra degree of freedom provided by fluid antenna (FA), incorporating FA into IDET systems presents a promising approach to enhance energy efficiency performance. This paper investigates a FA assisted IDET system, where the transmitter is equipped with multiple FAs and transmits wireless signals to the data receiver (DR) and the energy receiver (ER), both of which are equipped with a single traditional antenna. The switching delay and energy consumption induced by port selection are taken into account in IDET system for the first time. We aim to obtain the optimal beamforming vector and the port selection strategy at the transmitter, in order to maximize the short-term and long-term WET efficiency, respectively. The instant sub-optimal solution is obtained by alternatively optimizing the beamforming vector and port selection in each transmission frame, while a novel constrained soft actor critic (C-SAC) algorithm is proposed to find the feasible policy of port selection from the long-term perspective. Simulation results demonstrate that our scheme is able to achieve greater gain in terms of both the short-term and long-term WET efficiency compared to other benchmarks, while not degrading WDT performance.
@@ I-A Background
Physical design of FA mainly focuses on liquid-based antennas, reconfigurable pixel-based antennas and stepper motor based antennas [4]. The liquid based FA consists of liquid radiating elements (e.g., mercury, ga, galinstan, etc.) enclosed within a dielectric holder and controlled by a dedicated microelectromechanical system (MEMS) [5]. The reconfigurable pixel based FA is composed of an array of metasurface elements or reconfigurable sub-wavelength structures that can dynamically modify their electromagnetic properties [6][7] and the stepper motor based FA is known as the movable antenna (MA)[8]. Compared to the high hardware costs of pixel-based antennas, which stem from a large number of fixed antennas, liquid-based FAs and MAs are more cost-effective as they require fewer antenna elements and RF chains. However, liquid-based FAs and MAs incur delay and energy consumption during the movement, which potentially limiting their use in some communication scenarios. Therefore, evaluating the impact of FA movement is crucial for achieving higher energy efficiency and spectral efficiency of IDET systems.
@@ I-B Related works
In order to establish an energy-efficient interconnected network, IDET leverages RF signals to supply energy for low-power wireless devices. Plenty of works have been dedicated to the transceiver design of IDET system, aiming to improve the performance trade-off between WDT and WET. For instance, Kwon et al. [9] developed a transceiver design for a mmWave IDET system, including the transmit power allocation, beamforming and receive power splitting. Zhang et al. [10] considered a practical IDET scenario, where the data receivers (DRs) and energy receivers (ERs) were located in the far- and near-field regions of the extremely large-scale array base station (BS), respectively. The weight sum-power harvested at ERs was maximized by jointly optimizing the beamforming vectors and transmit power under the constraints of maximum sum-rate. With the assistance of reconfigurable intelligent surface (RIS) [11, 12, 13], the performance of IDET system can be further improved. Wu et al. [11] studied a sets of RIS assisted IDET system, where the beamforming vectors and reflect phase shifts were jointly optimized to minimize the transmit power, subject to the quality-of-service (QoS) constraints for all users. Yaswanth et al. [12] jointly optimized the beamforming matrices at both base station and RIS to minimize the transmit power while satisfying the minimum DR and ER constraints, thereby contributing to an energy efficient transmission scheme. More recently, with the emergence of integrated sensing and communication (ISAC) [16, 14, 15], efforts have been made to integrate ISAC into IDET systems, offering further opportunities for energy-efficient and multifunctional network designs.
@@ I-C Motivation and Contributions
Although previous works, including our own work [26] have shown that deploying FAs in IDET systems can achieve outstanding performance than traditional antennas, they have not considered the impact of switching delay and energy consumption associated with antenna movement, which may decrease the throughput and energy efficiency of IDET system. In IDET system, the WDT performance and WET performance can be improved by moving FAs into strategic positions, where the signal strength is maximized, or interference is minimized, or a trade-off between both factors is achieved. However, the movement of FAs introduces both delay and energy consumption, which reduces the effective transmission time within each frame, thereby lowering the achievable throughput of DR and harvested energy of ER. Therefore, it is crucial to balance the channel gain brought from the FAs and the energy efficiency loss caused by their movement. Motivated by this, we investigate the impact of delay and energy consumption due to FA movement on the IDET performance from both short-term and long-term perspectives. To improve the energy efficiency of FA-assisted IDET system, we optimize beamforming design and port selection to maximize the WET efficiency at the ER. However, this energy-oriented design may misalign with the optimal direction for the DR, potentially compromising WDT performance. To address this trade-off, we incorporate a minimum throughput constraint at the DR in both short-term and long-term optimization, ensuring reliable communication while enhancing WET efficiency. The contributions are summarized as follows:
@@
  • The WET efficiency is maximized by optimizing the beamforming vector and port selection from both short- and long-term perspective, while ensuring the harvested energy constraint and achievable throughput constraint. In short-term optimization, we transform the original fraction problem into linear problem by using Dinkebach’s algorithm, while the port selection and beamforming vector are then optimized alternatively by using feasible point pursuit-successive convex approximation (FPP-SCA) and semidefinite relaxation (SDR).
@@
Notations: In this paper, boldface upper letters, boldface lower letters, and lower letters denote matrices, vectors, and scalars, respectively. $\mathbb{C}^{x\times y}$ and $\mathbb{R}^{x\times y}$ represent the set of complex matrices and real matrices with the dimension of $x\times y$, respectively. $|x|$ represents the amplitude of a complex number $x$. $\boldsymbol{x}||$ and $\boldsymbol{x}||_{1}$ denote the 2-norm and 1-norm of a vector $\boldsymbol{x}$, respectively. The notation $(\cdot)^{T}$ and $(\cdot)^{H}$ refer to the transpose and the conjugate transpose of a vector or matrix, while $\mathbf{E}$ represents the expectation operator.
@@ II System Model
As shown in Fig. 1, we consider a multiple FAs assisted IDET system, which is comprised of a transmitter, a DR and an ER. The transmitter is equipped with $N_{t}$ FAs, each of which has the same size and parameters without loss of generality. The DR and ER are both equipped with a single traditional antenna due to the limited hardware space. We assume that the transmit antennas are sufficiently spaced (larger than $\lambda/2$ or equal to $\lambda/2$ ), such that their spatial correlation and mutual coupling can be considered negligible [20]. Consequently, the channels of distinct fluid antennas are independent, consistent with the assumptions commonly used in classic MIMO systems. However, the channels among different ports in the same antenna are strongly spatially correlated due to their close physical proximity. The DR and ER are randomly distributed in two different directions of the transmitter, which allows the beamforming vector to be optimized with sufficient spatial flexibility. Moreover, considering that the ER typically requires a higher harvested energy to meet its operational threshold, it is generally deployed closer to the transmitter to compensate for the severe pathloss in wireless energy transfer. The set of FAs and FA’s port are denoted by $\mathcal{N}_{t}=\left\{1,\cdots,N_{t}\right\}$ and $\mathcal{N}=\left\{1,\cdots,N\right\}$, respectively.
@@ II-B Wireless channel model
On the other hand, the channels between the transmitter and receivers are time-varying and exhibit temporal correlation across transmission frames. A widely accepted method to model channel evolution is to adopt a first-order Gaussian-Markov process with a one-step correlation coefficient derived from Jakes’ model [30, 31]. Specifically, the first-order Gaussian-Markov channel model can be expressed as
$\boldsymbol{h}_{i}(t)=\rho\boldsymbol{h}_{i}(t-1)+\sqrt{\frac{1-\rho^{2}}{PL_{h,i}}}\boldsymbol{\delta}(t),$
$\boldsymbol{g}_{i}(t)=\rho\boldsymbol{g}_{i}(t-1)+\sqrt{\frac{1-\rho^{2}}{PL_{g,i}}}\boldsymbol{\delta}(t)$ (5)
where $\rho\in[0,1]$ is the temporal correlation coefficient and $\boldsymbol{\delta}\sim\mathcal{CN}(0,\boldsymbol{I}_{N})$ is the random difference with unit-variance complex Gaussian i.i.d. in time.
@@ II-C Signal model
where $\eta\in(0,1]$ represents the energy conversion efficiency [^1] In general, considering a nonlinear energy harvest model is more practical, but the received power at ER is much small due to the signal attenuation, enabling the EH circuit works in linear region of the nonlinear model [32]. Thus the linear EH model can be used to approximate the nonlinear model.. Assuming that the power actuating the liquid metal moving is a constant as $P_{u}$, the total energy consumption in coherence time is expressed as
[…]
@@
[…]
In this paper, we aim to investigate both short-term and long-term WET efficiency maximization problem for the FA assisted IDET system, subject to the throughput constraint of DR. Specifically:
@@
  • Conversely, the long-term optimization framework adopts a holistic perspective by considering temporal correlations of channels and state evolution across multiple transmission frames, which can achieve more stable long-term WET efficiency. Nevertheless, the long-term optimization based DRL algorithm requires prior experience and the training is time-consuming.
Note that when port switching delay and energy consumption are explicitly incorporated into the WET efficiency optimization, the system must balance the channel gain brought by port reconfigurability against the efficiency loss caused by antenna movement. This trade-off substantially increases the problem complexity and highlights the need for dedicated cost-aware optimization frameworks.
@@ III-A Problem Formulation
In this section, we optimize the beamforming vector and port selection to maximize the WET efficiency in each transmission frame by only exploiting the current channel state information (CSI). For the sake of simplicity, we omit the time index $t$ in the expression. Therefore, the optimization problem can be formulated as
$\text{(P1)}:\max_{\boldsymbol{w},\{k_{i}\}}\quad$ $\eta_{EE},$ (13)
@@
where $\boldsymbol{G}=\mathcal{M}\left\{\boldsymbol{g}_{i},i\in\mathcal{N}_{t}\right\}$ and $\boldsymbol{D}=\mathcal{M}\left\{\boldsymbol{d}_{i},i\in\mathcal{N}_{t}\right\}$. Note that the objective function is a complicated fraction and the beamforming vector and the port activation indicators are intricately coupled, thus (P2) is non-convex optimization problem and is hard to solve directly. Referring to previous work [33], Dinkelbach’s algorithm is widely applied to transform fractional objective function into linear function by introducing an auxiliary variable $q$, making the optimization problem more tractable. Note that when the single-ratio problem is a concave-convex fractional programming, the Dinkelbach’s algorithm can converge to the global optimal solutions. Therefore, we aim to reformulate the objective function into a concave-minus-convex form in the subsequent transformation. By relaxing the binary constraints and denoting $\boldsymbol{C}=\mathcal{M}\left\{\boldsymbol{1}_{i}^{N},i\in\mathcal{N}_{t}\right\}$, (P2) can be reformulated as
$\text{(P3)}:\max_{\boldsymbol{w},\hat{\boldsymbol{a}},\tau_{s}}\quad$ $E_{h}-qE_{c},$ (17)
@@
$\hat{\boldsymbol{a}}\in[0,1]^{NN_{t}}.$ (17g)
In (P3), $q$ is iteratively updated by
$q^{n}=\frac{E_{h}^{n}}{E_{c}^{n}}$ (18)
where $n$ is the iteration index. For a fixed parameter $q$, (P3) is still a non-convex problem, thus we develop a AO algorithm to tackle it by solving $\hat{\boldsymbol{a}}$, $\tau_{s}$ and $\boldsymbol{w}$ iteratively.
@@ III-B Port selection optimization
[…]
@@
(P4) contains non-convex terms $|\hat{\boldsymbol{a}}^{T}\boldsymbol{G}\boldsymbol{w}|^{2}$ and $|\hat{\boldsymbol{a}}^{T}\boldsymbol{H}\boldsymbol{w}|^{2}$ in both the objective function (20) and constraints (20a)-(20b), while maintaining convexity in all other terms with respect to optimization variables $\hat{\boldsymbol{a}}$ and $\tau_{s}$. Although semidefinite relaxation (SDR) approach can be used to linearize the non-convex components, its application would induce significant dimensionality expansion of the optimization variables, rendering it computationally prohibitive for FA-assisted IDET system with large port number $N$. The FPP-SCA algorithm [34] is proposed for approximately solving general quadratic constraint quadratic problems (QCQPs), while a feasible point is guaranteed by adding slack variables. For any $\hat{\boldsymbol{a}}$$\hat{\boldsymbol{x}}$, we have $(\hat{\boldsymbol{a}}-\hat{\boldsymbol{x}})^{T}\boldsymbol{S}(\hat{\boldsymbol{a}}-\hat{\boldsymbol{x}})\geq 0$, where $\boldsymbol{S}=\eta\boldsymbol{G}\boldsymbol{w}\boldsymbol{w}^{H}\boldsymbol{G}^{H}$. Expanding the right-hand side of the inequality, we obtain
@@
$\text{(P5)}:\max_{\hat{\boldsymbol{a}},\tau_{s},s_{1},s_{2}}$ $2Re\left\{\hat{\boldsymbol{a}}^{T}\boldsymbol{S}\hat{\boldsymbol{x}}\right\}-\hat{\boldsymbol{x}}^{T}\boldsymbol{S}\hat{\boldsymbol{x}}-p(s_{1}+s_{2})$
@@
$s.t.\ \$$2Re\left\{\hat{\boldsymbol{a}}^{T}\boldsymbol{S}\hat{\boldsymbol{x}}\right\}-\hat{\boldsymbol{x}}^{T}\boldsymbol{S}\hat{\boldsymbol{x}}+s_{1}\geq\frac{E_{th}}{T-\tau_{s}},$ (22a)
@@
where $p$ is slack penalty, $s_{1}$ and $s_{2}$ are slack variables, $\boldsymbol{F}=\boldsymbol{H}\boldsymbol{w}\boldsymbol{w}^{H}\boldsymbol{H}^{H}$. Given an initial point $\hat{\boldsymbol{x}}$, (P5) is a convex optimization problem and can be optimally solved with the assistance of convex optimization tools. Due to the application of first-order Taylor expansion, the solution obtained from (P5) is the sub-optimal of (P4), and the objective value of (P5) provides a lower bound for (P4).